OpenStudy (anonymous):

The length of a rectangle is three times its width. If the area of the rectangle is 75in^2 , find its perimeter.

OpenStudy (nurali):

Let width of the rectangle is x => Length is 3x Now Area = 3x.x = 3x^2 3x^2 = 75 => x^2 = 75/3 => x^2 = 25 => x = +/- 5 But x = -5 is not possible because side can't be negetive. hence, Width = 5 Length = 3 x 5 = 15 Perimeter = 2 (width + Length) = 2 (5 + 15) = 2 x 20 = 40

OpenStudy (anonymous):

So, you know that length one is a, and length two is 3a. That means that the area of the square is 3a^2. Setting this equal to 75, we find that a is equal to sort(25) which is equal to five. Finding the perimiter is simple, as it is 3a + a + 3a + a = 8a = 40

OpenStudy (anonymous):

i the peremiter 2 or 40?

OpenStudy (nurali):

Perimeter = 2 (width + Length) = 2 (5 + 15) = 2 x 20 = 40

OpenStudy (nurali):

if correct please click best response.

OpenStudy (nincompoop):

what is 15 x 5?